OpenStudy (anonymous):

How to simplify 11/5 (ln2 - 1/2 (ln9) + 1/2 (ln6)?

OpenStudy (anonymous):

You have no variables, so can't you just solve it?

OpenStudy (anonymous):

i need to be in exact form

OpenStudy (anonymous):

Oh, right. Sorry, wasn't thinking.

OpenStudy (anonymous):

i dont know if i can just bring the 1/2 to the 9 , so it will be 3

OpenStudy (anonymous):

Well, \(1/2 \ln 9\) is \(\ln 3\), right?

OpenStudy (anonymous):

Yeah, because \(a \log x = \log x^a\)

OpenStudy (anonymous):

then will the final answer be (ln 2sqrt 6/ 3)?

OpenStudy (anonymous):

\[\frac{11}{5} (\ln 2 - \ln 3 + \ln \sqrt{6}) = \frac{11}{5} (\ln \frac{2 \sqrt{6}}{3}) \text{ or } \ln (\frac{2 \sqrt{6}}{3})^{\frac{11}{5}}\] Is that math right?